2. Examine the rise in gasoline prices from 1997 to 2006. The price of regular unleaded gasoline in January 1997 was $1.26, and in January 2006, the price of regular unleaded gasoline was $2.31 (“Consumer price index,” 2006). Use the coordinates (1997, 1.26) and (2006, 2.31) to find the slope, or rate of change, between the two points. Describe how you arrived at your answer.
Answer by stanbon(75887)

Use the coordinates (1997, 1.26) and (2006, 2.31) to find the slope, or rate of change, between the two points.
---------------
slope = (2.31-1.26)/(9) = 0.1167
---
Form: y = mx+b
1.26 = 0.1167*1997 + b
-----
b = 1.26-0.1167*1997
-----
b = -231.79
=================
Equation:
y = 0.1167x - 231.79
===========
Cheers,
Stan H.